### v1

# JWST’s little red dots: an emerging population of young, low-mass AGN cocooned in dense ionized gas

Abstract
JWST has uncovered large numbers of compact galaxies at high redshift with broad hydrogen/helium lines. These include the enigmatic population known as “little red dots” (LRDs)[1, 2]. Their nature is debated, but they are thought to be powered by supermassive black holes (SMBHs)[3, 4] or intense star formation[5]. They exhibit unusual properties for SMBHs, such as black holes that are overmassive for their host galaxies[4] and extremely weak X-ray[6, 7, 8, 9] and radio[10, 11, 12] emission. Using the highest-quality JWST spectra, we show here that the lines are broadened by electron scattering with a narrow intrinsic line core. The data require high electron column densities and compact sizes (light days), which, when coupled with their high luminosities can only be explained by SMBH accretion. The narrow intrinsic cores of the lines imply upper limits on the black hole masses of M⊙, two orders of magnitude lower than previous estimates. These are among the lowest mass SMBHs known at high redshift and suggest that this is a population of young, rapidly growing SMBHs. They are enshrouded in a dense cocoon of ionized gas, probably related to their youth, from which they are accreting close to the Eddington limit. Reprocessed nebular emission from the dense cocoon dominates the optical spectrum, explaining most LRD spectral characteristics and helping to suppress radio and X-ray emission.[13, 14]
Two early clues suggested that these galaxies may be affected by Compton-thick ionized gas: Balmer absorption features are often observed in the broad lines[2, 15, 16], and if they are active galactic nuclei (AGN), their X-rays might be suppressed by photoelectric absorption in Compton-thick ionized gas in the broad-line region (BLR) near the SMBH[6]. These ideas inspired us to investigate the H line profiles to look for electron scattering signatures produced by the large ionized gas column densities.
To test different broad line shapes, we constructed a sample of all broad line galaxies (H km s-1) with high signal-to-noise ratio (SNR) JWST/NIRSpec medium-resolution () spectra [17] in the DAWN JWST Archive (DJA)[18]. Details of the sample and selection criteria are presented in the Methods and Extended Data Table 2. Our search yielded 12 objects at and 18 additional objects for a combined ‘stacked’ spectrum at . High-resolution () data were used where available (Objects A and C) and gave consistent results but better constrained narrow component widths.
Our object selection is strongly linked to LRDs, many of which show broad lines. LRDs are spatially compact with characteristic ‘v’-shaped optical-UV spectra[1, 19], due to a change of slope close to the Balmer limit wavelength, nm[20]. Despite our selection being based only on H linewidth and SNR criteria, our systems are all spatially very compact (Fig. 2), and often have a clear slope change at the Balmer limit in the restframe spectra (see Extended Data Fig. 1).
Electron scattering in a dense ionized gas produces lines with exponential profiles[21, 22, 23], while Gaussian or centrally broad profiles are expected from the typical Doppler broadening due to bulk gas motion around SMBHs[24]. We therefore compare two basic models for the broad H component: a Gaussian and a double-sided exponential. Raman scattering[9] or turbulence[24] give rise to broad Lorentzian lines. However, since we find the exponential to be a better or equivalent fit in all cases, for clarity of presentation, we discuss the Lorentzian model in the Methods section. In addition to the broad component and a local continuum in the H region, we also model the narrow emission lines that cannot easily be masked out: H and the [N ii] lines 6549,6585. Although rare in star-forming galaxies or AGN at lower redshift, relatively narrow Balmer absorption features are also a common feature in these objects[2] and can be seen on top of the H line of many of our sample objects (Fig. 3). They are accounted for in our fits as an absorption line in object E and as P Cygni absorption/emission features in A and D. For the Gaussian/exponential comparison, we examine the quality of these simpler fits and find that while the Gaussian fit leaves systematic residuals with a characteristic ‘W’-shape around the central region (Fig. 1), the exponential model is superior in almost every case. The significance of the fit improvement increases with the SNR of the spectra (Extended Data Fig. 6), an indication that the exponential model is the better fit in the general case, even where the spectra do not have high enough SNR to discriminate. The comparison fits to all spectra and the stacked spectrum are shown in Extended Data Fig. 4. Fit statistics are given in Extended Data Table 3. To illustrate more explicitly the exponential nature of the lines, we plot the data with a logarithmic ordinate scale (Fig. 1, right and Extended Data Fig. 5). The lines are symmetric and trace a straight line on a semi-logarithmic plot over several orders of magnitude, consistent with moderate optical depth electron-scattering[22]. These results imply that the primary line-broadening mechanism is electron-scattering through a Compton-thick medium and not primarily bulk Doppler motions, immediately excluding high mass SMBHs. The symmetry of the lines suggests that any outflow of the scattering medium is less than a few hundred km s-1 (ref. [23]).
After establishing that the basic broad line shape is dominated by electron-scattering effects, we allow that they also have an intrinsic Doppler line core. For our fiducial model we therefore use a Gaussian convolved with an exponential, with the widths free to vary, to measure the intrinsic width of the Doppler lines. Fits for all objects are shown in (Fig. 3). In most cases the intrinsic line width is small; we find average intrinsic velocities of approximately 300 km s-1 for ten of the twelve objects in the sample. For objects B and G we recover relatively large Gaussian intrinsic widths, about 2000 km s-1. However, we caution that specifically these spectra have the most complex line shapes in our sample, with three apparent peaks. It is probable that our simple model is not adequately reproducing complex radiative transfer effects through a (highly) optically thick medium. We therefore reserve judgement on whether objects B and G really have such large Doppler velocities.
The characteristic properties of the gas can be inferred from the line widths. First, we infer the electron column density, , from the width of the lines[23, 22], assuming an electron temperature between 10,000 and 30,000 K. We use our own Monte Carlo electron scattering code with a simple shell geometry to calculate the relationship between the electron-scattering optical depth, , the temperature, and the characteristic width of the exponential (see Methods). We find – 2.8, implying – cm-2 for the sample (see Table 1 and Extended Data Fig. 8). The absence of strong broadening of the [O iii] doublet, while the H line next to it is broadened, suggests that the electron volume density, , in the scattering region, must be at least several times the critical density for these lines. The characteristic spherical size of the scattering region, , can be inferred by taking the ratio of the column and volume densities in a simple constant density sphere, i.e. . Assuming , we find sizes of at most a few hundred light days and ionized gas masses of order M⊙. If we assume densities more in line with estimates in the literature for these objects (i.e. [26, 27], the sizes would be equivalently smaller, about a few light days.
The dense Compton-thick gas inferred from the line wings offers a natural explanation for the Balmer absorption lines frequently seen in the broad-line systems at high redshift[2, 27]. Given the extreme column and volume densities of the ionised gas inferred here, the fraction of the gas in the level should provide sufficient optical depth to produce these features. We model the features with P Cygni profiles in the spectra of A and D. The P Cygni model gives a much better fit than a pure absorption line due to the sharpness of the transition between the emission peak and the blueshifted absorption feature, characteristic of the P Cygni profile[28]. The properties of the absorption features indicate a largely spherical gas distribution, with mild outflow velocities below a few hundred km s-1. In cases of high optical depth and lower velocities, radiative transfer effects may instead induce a complex self-absorption feature in the line centre. The presence of narrow emission lines from the rest of the galaxy may make the profile even more complicated or conceal these features, making it difficult to assess how common such Balmer absorption is. A full investigation of the Balmer absorption lines is beyond the scope of this paper and we defer their analysis to future work.
We now turn to the origin of the lines in these systems. The line widths in most of our sample could be explained by very mild outflows, consistent with the photospheric velocities we find for the P Cygni lines in objects A and D (about 200–300 km s-1). Such a mild outflow could be due to feedback from a burst of star formation. The largest maximum radii we infer are roughly consistent with the smaller star clusters observed in R136 in the LMC and NGC 253[29] in the local universe, or in the Sunrise Arc at or the Cosmic Gems system at redshift [30]. However, the luminosity densities of those star clusters are typically about erg s-1 per pc2 or per pc3. The more luminous systems in our sample, on the other hand, have ionizing luminosities of erg s-1, based on observed H luminosities around erg s-1 [31, 32]. This is nearly four orders of magnitude higher than the highest luminosity density star-formation known. Distributing the inferred luminosity across multiple sources over a larger area or volume (e.g. many accreting extremely massive stars[33]) doesn’t resolve this problem because the inferred gas masses require luminosities L⊙ to keep them ionized. Because these inferred luminosity densities are based on the profiles of the lines, which represent the average conditions of the emitting regions, this requires each of the putative individual sources to be this luminous. These considerations essentially rule out a star-forming origin for the broad lines. Production of such large ionising luminosities in regions significantly smaller than a few hundred light days can only realistically be powered by accretion onto a SMBH, i.e. an active galactic nucleus (AGN). The AGN radius-luminosity relation implies a H broad-line region radius[34] of approximately five light days for a H luminosity of erg s-1. This is consistent with the upper bounds inferred above. If this radius is accurate, it suggests that the volume density in the scattering regions may be somewhat lower than typical BLR cloud densities.
As mentioned above, the AGN interpretation of the broad line objects and LRDs discovered by JWST faces several challenges: the large inferred black hole masses, both compared to their host galaxies, and as a function of the age of the universe and their cosmic number density, and the lack of X-ray and radio detections. First, as the electron scattering dominates the line widths in these systems, we need to use the intrinsic line widths prior to scattering, to infer the black hole masses, assuming these intrinsic line widths are entirely due to Doppler orbital motions around the SMBH. These are an order of magnitude lower than inferred with a simple Gaussian fit to the full line profile. The inferred black hole masses, using standard single-epoch black hole mass scaling relations[35], are therefore lower by about a factor of a hundred (Table 1). Our estimates of the maximum SMBH masses are consistent with the typical host galaxy stellar mass found at lower redshifts[35], as shown in Fig. 4, alleviating much of this problem. The much lower black hole masses also alleviate the number density and early formation and growth problems. Second, Compton-thick ionized gas will hinder X-ray[26] and suppress radio[27] emission, by photoelectric and free-free absorption[36] respectively. This is true in X-rays even at low metallicity due to He absorption[37]. The ionized column densities we infer here will attenuate the X-rays by only a factor of a few (see Methods)–less than the dex required to satisfy the observations[6, 7, 8]. Steep hard X-ray spectral slopes and/or relatively low energy power-law cut-offs are also required to suppress the X-rays[6], both of which are found in narrow-line Seyfert 1 objects[38]. Narrow-line Seyfert 1s are believed to be low-mass, high accretion rate SMBHs[39].
The high density gas reprocesses essentially all of the Lyman continuum radiation from the central source, resulting in extremely high Lyman optical depths that cause the self-reversal seen as absorption in the line centres of some of the Balmer lines. The bulk of the ionizing flux from the AGN is therefore emitted in the Balmer and Paschen regions of the spectrum, implying that nebular gas emission must be a major component of the optical and NIR spectra of these sources, giving rise to Balmer, Paschen and He i emission lines and continua and the two-photon continua that would dictate the spectral shape[13, 40]. In addition, while many of these systems appear in blue star-forming galaxies based on the spatially-extended nebulosity surrounding the point source[41], many may be entirely dominated by the nebular emission at optical and even near UV wavelengths. We also note that our inferred electron-scattering optical depths are all (see Extended Data Fig. 8). This suggests three things. First, that there may be a population that is even more heavily obscured, suppressing line emission due to self-absorption, and giving rise to spectra that appear in absorption rather than emission, including a Balmer break instead of a jump. Second, it also suggests that the gas distribution is close to spherical without a large opening, because scattered light will be preferentially emitted along the lowest column density sightlines. The fact that we are not detecting large numbers of low-mass, rapidly accreting AGN without these signatures indicates that LRDs are the primary population and that the gas cocoon in these SMBHs is a near-universal, young phase of their accretion-powered growth. Third, that the line ratios of the Balmer lines should be very different from the standard case B values due to very high Lyman optical depths [42], so that high H/H ratios do not imply dust extinction.
| ID | Field PID-SRCID | Redshift | MBH | M | M⋆ | ||
| ( cm-2) | ——— () ——— | ||||||
| A | GN 1181-68797 | 5.0405 | |||||
| B | GN 2674-14 | 5.1826 | |||||
| C | GN 1181-73488 | 4.1327 | |||||
| D | EGS 4233-42046 | 5.2757 | |||||
| E | EGS 4233-49140 | 6.6847 | |||||
| F | EGS 1345-1244 | 4.4771 | — | ||||
| G | EGS 4233-58237 | 3.6505 | |||||
| H | EGS 4106-51623 | 4.9511 | |||||
| I | GN 1181-53501 | 3.4294 | |||||
| J | GN 1181-38147 | 5.8694 | |||||
| K | EGS 4233-50052 | 5.2392 | |||||
| La | EGS 4233-60935 | 5.2877 | |||||
| Sb | — | — | — | — | |||
aObject L has two estimates of corresponding to two distinct FWHM solutions.
b for the stack is a median of the individual masses.
Another conundrum of these sources was that they are found extremely early in the history of the universe with (previously inferred) very large masses and low accretion rates[16]. If they really represent the early stages of black hole growth and require high accretion rates over very long times to grow rapidly[43], they should be found in high accretion states when they are brighter. The lower black hole masses we infer here solve this puzzle. We calculate a mean Eddington ratio around unity for our sources A–L, excluding objects B and G, assuming the H line is a few percent of the bolometric luminosity. These sources are indeed accreting close to, or in some cases, slightly above, the Eddington limit, as would be expected if we are discovering a population of early, fast-growing SMBHs. As our sample is likely to be biased in favour of more luminous systems, the bulk of the population may have even lower black hole masses, lower accretion rates, and/or greater dust extinction. This suggests that SMBH masses at the M⊙ level could be a substantial part of the population.
All of this points to a solution to JWST’s riddle of LRDs/compact broad-line objects. They are intrinsically narrow-line AGN, i.e. young[44], low-mass (– M⊙), Eddington-accreting SMBHs[39], buried in a thick cocoon of gas, presumably related to their early stage of formation[45]. Their high accretion rates produce copious UV emission that ionizes their gas cocoon and at the same time efficiently cool and weaken the corona, suppressing their hard X-rays[39]. The ionized gas hinders the escape of radio and X-rays, while reprocessing almost all the Lyman radiation into nebular optical emission, producing the broadened Balmer lines and continuum breaks that characterise the classic ‘v’-shape spectrum[20]. Many galaxies have been found with stellar masses and ages that are so massive and old that they are difficult to explain in standard scenarios[46, 47]. Our results and other recent work[13, 14] indicate that galaxies with red colours and even those with Balmer breaks, may, instead of being massive, quiescent or post-starburst galaxies, be low-mass AGN systems with extreme optical depth ionized gas. The distribution of electron column densities and the lack of a similar or larger population of low-mass, high accretion AGN without the nebular cocoon hints that we may be observing the bulk growth of SMBHs in the phase when they are surrounded by a quasi-spherical, dense gas shell, before angular momentum transport has flattened their accretion into a well-defined disk.
0.1 References

### v2

@@
 Their nature is debated, but they are thought to be powered by supermassive black holes (SMBHs)[3, 4] or intense star formation[5].
-They exhibit unusual properties for SMBHs, such as black holes that are overmassive for their host galaxies[4] and extremely weak X-ray[6, 7, 8, 9] and radio[10, 11, 12] emission.
+They exhibit unusual properties for SMBHs, such as black holes that are overmassive for their host galaxies[4] and extremely weak X-ray[6, 7, 8, 9, 10] and radio[6, 11, 12, 13] emission.
 Using the highest-quality JWST spectra, we show here that the lines are broadened by electron scattering with a narrow intrinsic line core.
@@
 These ideas inspired us to investigate the H line profiles to look for electron scattering signatures produced by the large ionized gas column densities.
-To test different broad line shapes, we constructed a sample of all broad line galaxies (H km s-1) with high signal-to-noise ratio (SNR) JWST/NIRSpec medium-resolution () spectra [17] in the DAWN JWST Archive (DJA)[18].
+To test different broad line shapes, we constructed a sample of all broad line galaxies (H km s-1) with high signal-to-noise ratio (SNR) JWST/NIRSpec medium-resolution () spectra [18] in the DAWN JWST Archive (DJA)[19, 20].
 Details of the sample and selection criteria are presented in the Methods and Extended Data Table 2.
@@
 This is true in X-rays even at low metallicity due to He absorption[39].
-The ionized column densities we infer here will attenuate the X-rays by only a factor of a few (see Methods)–less than the dex required to satisfy the observations[6, 7, 8].
+The ionized column densities we infer here will attenuate the X-rays by only a factor of a few (see Methods)–less than the dex required to satisfy the observations[6, 8, 7, 10].
 Steep hard X-ray spectral slopes and/or relatively low energy power-law cut-offs are also required to suppress the X-rays[10], both of which are found in narrow-line Seyfert 1 objects[40].
@@
 This suggests three things.
-First, that there may be a population that is even more heavily obscured, suppressing line emission due to self-absorption, and giving rise to spectra that appear in absorption rather than emission, including a Balmer break instead of a jump.
+First, that there may be a population that is even more heavily obscured, suppressing line emission due to self-absorption, and giving rise to spectra that appear in absorption rather than emission, including a Balmer break instead of a jump[45, 46].
 Second, it also suggests that the gas distribution is close to spherical without a large opening, because scattered light will be preferentially emitted along the lowest column density sightlines.
@@
 Third, that the line ratios of the Balmer lines should be very different from the standard case B values due to very high Lyman optical depths [47], so that high H/H ratios do not imply dust extinction.
-| ID | Field PID-SRCID | Redshift | MBH | M | M⋆ | ||
+| ID | Survey-Field-MSAID | Redshift | MBH | M | M⋆ | ||
 | ( cm-2) | ——— () ——— | ||||||
-| A | GN 1181-68797 | 5.0405 | |||||
-| B | GN 2674-14 | 5.1826 | |||||
-| C | GN 1181-73488 | 4.1327 | |||||
-| D | EGS 4233-42046 | 5.2757 | |||||
-| E | EGS 4233-49140 | 6.6847 | |||||
-| F | EGS 1345-1244 | 4.4771 | — | ||||
-| G | EGS 4233-58237 | 3.6505 | |||||
-| H | EGS 4106-51623 | 4.9511 | |||||
-| I | GN 1181-53501 | 3.4294 | |||||
-| J | GN 1181-38147 | 5.8694 | |||||
-| K | EGS 4233-50052 | 5.2392 | |||||
-| La | EGS 4233-60935 | 5.2877 | |||||
+| A | JADES-GN-68797 | 5.0405 | |||||
+| B | 2674-GN-14 | 5.1826 | |||||
+| C | JADES-GN-73488 | 4.1327 | |||||
+| D | RUBIES-EGS-42046 | 5.2757 | |||||
+| E | RUBIES-EGS-49140 | 6.6847 | |||||
+| F | CEERS-EGS-1244 | 4.4771 | — | ||||
+| G | RUBIES-EGS-58237 | 3.6505 | |||||
+| H | 4106-EGS-51623 | 4.9511 | |||||
+| I | JADES-GN-53501 | 3.4294 | |||||
+| J | JADES-GN-38147 | 5.8694 | |||||
+| K | RUBIES-EGS-50052 | 5.2392 | |||||
+| La | RUBIES-EGS-60935 | 5.2877 | |||||
 | Sb | — | — | — | — | |||
@@
 The ionized gas hinders the escape of radio and X-rays, while reprocessing almost all the Lyman radiation into nebular optical emission, producing the broadened Balmer lines and continuum breaks that characterise the classic ‘v’-shape spectrum[22].
-Many galaxies have been found with stellar masses and ages that are so massive and old that they are difficult to explain in standard scenarios[46, 47].
+Many galaxies have been found with stellar masses and ages that are so massive and old that they are difficult to explain in standard scenarios[51, 6, 52].
 Our results and other recent work[14, 15] indicate that galaxies with red colours and even those with Balmer breaks, may, instead of being massive, quiescent or post-starburst galaxies, be low-mass AGN systems with extreme optical depth ionized gas.